A wire loop is 50mm in diameter and is oriented with its plane perpendicular to a magnetic field. How rapidly should the field change if a potential difference of 1.0V is to appear across the ends of the loop?


1
Expert's answer
2021-06-10T10:22:24-0400

According to Faraday's law, the induced electromotive force is


"E=\frac{\Delta \Phi}{\Delta t}=\frac{A·\Delta B}{\Delta t},\\\space\\
\frac{\Delta B}{\Delta t}=\frac EA=\frac{4E}{\pi d^2}=509.3\text{ T/s}."